# DBellQuant: Breaking the Bell with Double-Bell Transformation for LLMs Post Training Binarization

### v1

@@ Abstract
Large language models (LLMs) demonstrate remarkable performance but face substantial computational and memory challenges that limit their practical deployment. Quantization has emerged as a promising solution; however, its effectiveness is often limited by quantization errors arising from weight distributions that are not quantization-friendly and the presence of activation outliers. To address these challenges, we introduce DBellQuant, an innovative post-training quantization (PTQ) framework that achieves nearly 1-bit weight compression and 6-bit activation quantization with minimal performance degradation. DBellQuant uses Learnable Transformation for Dual-Bell (LTDB) algorithm, which transforms single-bell weight distributions into dual-bell forms to reduce binarization errors and applies inverse transformations to smooth activations. DBellQuant sets a new state-of-the-art by preserving superior model performance under aggressive weight and activation quantization. For example, on the Wikitext2 dataset, DBellQuant achieves a perplexity of 14.39 on LLaMA2-13B with 6-bit activation quantization, significantly outperforming BiLLM’s 21.35 without activation quantization, underscoring its potential in compressing LLMs for real-world applications.
@@ 1 Introduction
In recent years, the rapid advancement of large language models (LLMs) demonstrate exceptional performance in a variety of complex tasks that involve natural language understanding and generation achiam2023gpt ; dubey2024llama . However, these models often comprise hundreds of billions of parameters, posing significant challenges for their deployment in real-world applications because of the substantial computational and memory requirements (e.g. a 70B model requires around 150GB GPU memory), resulting in huge operational costs and unacceptable inference latency.
In this context, quantization, as an effective model compression technique, has garnered significant attention. Due to the sparsity of information density in the weights of LLMs wang2024bitnet ; yu2024super , weight binarization has emerged as a promising quantization scheme. Methods such as PB-LLM shang2023pb and BiLLM huang2024billm have significantly reduced the errors introduced by binarization by applying finer-grained processing to critical weights. Despite its theoretical advantages, the application of quantization in compressing LLMs continues to face numerous challenges egashira2024exploiting . For instance, systematic outliers in activation values can lead to substantial quantization errors, negatively impacting model accuracy sun2024massive . Although recent studies xiao2023smoothquant ; shao2023omniquant have proposed mitigating such errors by redistributing the scaling factors between weights and activation values. Other research directions ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm focus on leveraging Hadamard transformations to effectively eliminate outlier features. However, no binarization methods have taken care of simultaneous smoothing activation outliers. Consequently, the demand of higher bit-width activations during computation leads to significant computational overhead. BitNet a4.8 wang2024bitnet addresses this issue by employing quantization-aware training (QAT) to achieve a 1-bit LLM with 4-bit activations. Nevertheless, the QAT approach requires extensive computational resources and prolonged training time. Therefore, developing post-training quantization (PTQ) methods for binary compression that minimize performance loss while simultaneously smoothing activation and thus reducing activation bit-width remains a critical and challenging task.
@@
Figure 1: Performance on Wikitext2 dataset. DBellQunat outperforms weight-only quantization method under 8-bit activation setting.
To this end, we begin by revisiting the distribution characteristics of activations and weights in LLMs (Fig. 2(a)). We observe that the unimodal nature of weight distributions leads to substantial quantization errors, particularly in the case of low 1-bit quantization. Ideally, a dual-bell-shaped weight distribution (Fig. 2(b)) can effectively reduce binarization errors. This motivates a key question: Is it possible to transform weights into a dual-bell shape distribution while simultaneously addressing activation outliers to facilitate both activation quantization and weight binarization?
Building on these observations, we propose DBellQuant, a novel weight-activation quantization framework for efficient post-training quantization (PTQ). DBellQuant enables activation quantization while achieving near 1-bit weight compression with minimal accuracy loss. At the core of DBellQuant is a learnable transformation matrix that maps the weight distribution into a dual-bell form, while its inverse is applied to the input activations to smooth them. We begin by analyzing the characteristics of weight distributions suitable to binarization and theoretically derive the feasibility of applying a dual-bell transformation. Based on this, we then develop an efficient and lightweight algorithm, Learnable Transformation for Dual-Bell (LTDB). LTDB initializes the transformation matrix using an activation-aware strategy and optimizes it via a custom objective function that encourages the weights to cluster around two centers. This drives the formation of a symmetric dual-peak distribution, and an early stopping strategy is employed to ensure stable and efficient optimization. Through this equivalence-preserving transformation, weights originally exhibiting unimodal distributions– challenging for binarization– are mapped into near-symmetric dual-bell distributions (Fig. 2(b)), significantly reducing binarization error. Simultaneously, the inverse of the learned transformation is applied to the input activations, effectively scaling down outliers and smoothing the distribution. This makes the activations more amenable to quantization without altering the model’s output (Fig. 2(b)).
Experimental results demonstrate that the equivalent transformation strategy of DBellQuant significantly reduces the loss caused by weight binarization. For the first time under PTQ conditions, it achieves near 1-bit weight compression while simultaneously compressing activations to 6 bits.Across various LLMs and evaluation metrics, DBellQuant consistently achieves state-of-the-art results. For instance, on the Wikitext2 benchmark, we achieves a perplexity of 14.39 on LLaMA2-13B using only 6-bit activations (Fig. 1), significantly surpassing the performance of BiLLM, a method that only quantizes weights, which achieves a perplexity of 21.35.
@@ 2 Related Work
The massive parameter size of LLMs poses significant challenges in terms of memory consumption and computational efficiency. Therefore, quantization is crucial to compress these models, reducing resource requirements while preserving performance for practical deployment. LLMs quantization have introduced a variety of innovative techniques to enhance efficiency while maintaining accuracy. Works like GPTQ frantar2022gptq and OBQ frantar2023optimalbraincompressionframework minimizes reconstruction error by adjusting the remaining unquantized parameters in the block to compensate for the accuracy loss caused by quantization. LLM.int8()dettmers2022gpt3 and ZeroQuantyao2022zeroquant improve quantization accuracy by introducing additional grouping labels for customized quantization blocks. Other works like SmoothQuantxiao2023smoothquant and OmniQuant shao2023omniquant addresses activation outliers by redistributing scaling factors between weights and activations, migrating the quantization difficulty from activation to weights. Additionally, recent approaches leverage Hadamard transformations to suppress activation outliers ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm , while incoherence processing has been proposed for effective low-bit quantization chee2024quip2bitquantizationlarge ; tseng2024quipbetterllmquantization . Collectively, these advancements demonstrate that quantization techniques can be successfully scaled to multi-billion-parameter models, achieving substantial reductions in memory consumption and inference latency without compromising model performance.
@@
Binary quantization, an extreme low-bit quantization technique that reduces model weights and activations to binary values (e.g., -1 and +1 or 0 and 1), has gained significant attention for its ability to drastically cut memory usage and computational complexity, making it ideal for resource-constrained devices and efficient deployment of large-scale models. However, applying binary quantization to LLMs presents substantial challenges due to their sensitivity to precision loss, particularly in attention mechanisms and large embedding layers. BinaryBERT bai2020binarybert explored binary quantization for BERT, proposing selective preservation of critical weights in higher precision to mitigate performance degradation. In another direction, PB-LLM shang2023pb introduced a partially-binarized approach for LLMs, retaining a small fraction of salient weights in higher precision while binarizing the rest, enabling extreme low-bit quantization without sacrificing linguistic reasoning capabilities. Recent advancements include structural binarization techniques that leverage novel sparsity forms and standardized importance metrics to selectively binarize and sparsify LLM weights dong2024stbllmbreaking1bitbarrier , as well as strategies like alternating refined binarization and column-group bitmap methods to effectively reduce quantization error and address column deviations li2024arbllmalternatingrefinedbinarizations . These innovations collectively advance the feasibility of binary quantization for LLMs, pushing the boundaries of efficiency without compromising performance.
@@ 3.1 Binarization-Friendly Weight Redistribution
$\small\beta=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{W_{i,j}},\quad\boldsymbol{\widetilde{W}}=\text{Sign}(\boldsymbol{W}-\beta),\quad\alpha=\frac{1}{n}\sum_{i=1}^{n}|\boldsymbol{W_{i,j}}-\beta_{j}|$ (1)
$\text{Sign}(\boldsymbol{W_{i,j}})=\begin{cases}+1,&\text{if }\boldsymbol{W_{i,j}}>0,\\ -1,&\text{if }\boldsymbol{W_{i,j}}\leq 0,\end{cases},\quad\boldsymbol{W_{deq}}=\boldsymbol{\widetilde{W}}\cdot\alpha+\beta$ (2)
where $\beta$ is the shifting factor and $\alpha$ is the scaling factor for binarization. Previous studies huang2024slim have shown that neural network weights exhibit structured distributions along the channel dimension, with certain channels being more salient. The overall weight distribution typically follows a quasi-Gaussian pattern dettmers2022llmint88bitmatrixmultiplication , as does the channel-wise distribution (Fig. 6). Binarizing such weight matrices introduces significant quantization errors, which can severely degrade model performance.
In LLM binarization, a dual-bell distribution is theoretically more advantageous than a single-bell distribution due to its natural separation into two distinct clusters, which aligns well with binary quantization levels (e.g., -1 and 1), thereby minimizing quantization error. In contrast, single-bell distributions, concentrated around a single peak, often cause significant overlap when mapped to binary values, reducing representation accuracy (see Appendix A.4 for detailed analysis). However, LLM weight distributions typically exhibit single-bell characteristics, and conventional PTQ methods fail to effectively transform them for binarization. While QAT can reshape weight distributions into a dual-bell form through its learning objectives wang2023bitnetscaling1bittransformers , it requires substantial computational resources and prolonged training. To address this, we propose a more efficient PTQ method that rapidly converts single-bell distributions into dual-bell ones, optimizing binary quantization without the need for resource-intensive retraining.
@@ 3.2 Learnable Transformation for Dual-Bell Quantization
Learnable Transformation with Auxilary Matrix
@@
where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are parameters of the doubel-bell distribution. More detailed proof is shown in Appendix. A.3.
@@ 3.3 Dual-Transformation Optimizing Objectives
Our learning objective is to encourage all weight values to move closer to the two mean centers calculated by Eq. 2, ultimately forming a doubel-bell distribution with these two values as its peaks. During the binarization process, we denote these two points as $m_{1}$ and $m_{2}$ respectively. The simple way to set loss function is as follows:
@@
$\small\mathcal{L}_{\text{DTNP}}=\frac{\lambda_{DTNP}}{n}\sum_{i=1}^{n}\begin{cases}\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|}{\left|\text{m}_{1,i}\right|},&\text{if }\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|<\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|\\[10.0pt] \frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|}{\left|\text{m}_{2,i}\right|},&\text{otherwise.}\end{cases}$ (6)
@@ 3.4 Impact of the Inverse of Learnable Transformation Matrix on Activation Smoothing
Through the use of DTNP for training, we understand that the transformation matrix $T$ drives all values in the weights closer to the two mean centers calculated by Eq. 2. Prior research has shown that the distribution of weights tends to approximate a quasi-Gaussian distribution, with the majority of values being extremely small and close to zero, while only a very small fraction exhibit relatively large absolute values dettmers2022llmint88bitmatrixmultiplication . Theoretically, this implies that $T$ will contain many values greater than 1 to amplify the numerous near-zero absolute values, bringing them closer to the mean centers. At the same time, very few values of $T$ will be less than 1 to reduce the relatively rare large absolute values, aligning them similarly with the mean centers. In fact, when we visualize $T$ after training as shown in Fig. 5, we observe that less than 5% of its values are below 1. Consequently, the corresponding $T^{-1}$, which is multiplied with activations, has over 95% of its values below 1. This significantly reduces the quantization range of activations and suppresses the magnitude of outlier values within the activations. Visualization results can be seen in Appendix. A.6. As a result, this approach is particularly effective in further facilitating activation quantization.
@@ 3.5 Algorithm
The process of Learnable Transformation for Dual-Bell Transformation Algorithm is shown in Algorithm. 1, which adjusts the full-precision weight matrix $W$ using an activation-aware initialization transformation matrix $T$ over $N$ epochs. The algorithm iteratively minimizes dual-target loss functions to guide $W$ toward a dual-bell distribution while employing an early stopping mechanism based on the DTMD.
1:function LTDB($\boldsymbol{W},\mathbf{T},N$)
2:     Input: $\boldsymbol{W}\in\mathbb{R}^{n\times m}$ a full-precision weight matrix.
3:     Input: $\mathbf{T}\in\mathbb{R}^{1\times m}$ an activation-aware initialization transformation matrix.
4:     Input: $N$ the total number of epochs.
5:     Output: $\widetilde{\boldsymbol{W}}\in\mathbb{R}^{n\times m}$ the transformed weight matrix.
6:     for $\text{iter}=1,2,\dots,N$ do
7:         $\widetilde{\boldsymbol{W}}\leftarrow\mathbf{T}\odot\boldsymbol{W}$triangleright Perform element-wise multiplication
8:         $\mathcal{L}_{\text{DTMD}},\mathcal{L}_{\text{DTNP}}\leftarrow\text{LossFunc}(\widetilde{\boldsymbol{W}})$triangleright Compute the two dual-target losses for $\widetilde{\boldsymbol{W}}$
9:         $\mathcal{L}_{\text{DTNP}}.backward()$triangleright Use Dual-Target Normalized Proportional Loss for training
10:         if $\mathcal{L}_{\text{DTMD}}>\mathcal{L}_{\text{DTMD-Minimum}}$ then
11:              break $\triangleright$ Stop training based on Dual-Target Minimum Deviation Loss
12:         end if
13:         $\mathcal{L}_{\text{DTMD-Minimum}}\leftarrow\mathcal{L}_{\text{DTMD}}$
14:     end for
15:     return $\widetilde{\boldsymbol{W}}$
16:end function

### v2

@@ Abstract
Large language models (LLMs) demonstrate remarkable performance but face substantial computational and memory challenges that limit their practical deployment. Quantization has emerged as a promising solution; however, its effectiveness is often limited by quantization errors arising from weight distributions that are not quantization-friendly and the presence of activation outliers. To address these challenges, we introduce DBellQuant, an innovative post-training quantization (PTQ) framework that achieves nearly 1-bit weight compression and 6-bit activation quantization with minimal performance degradation. DBellQuant uses learnable transformation to map single-bell weight distribution to dual-bell distribution to reduce binarization error and smooth activations using inverse transformation. DBellQuant sets a new state-of-the-art by preserving superior model performance under aggressive weight and activation quantization. For example, on the Wikitext2 dataset, DBellQuant achieves a perplexity of 14.39 on LLaMA2-13B with nearly 1-bit weight and 6-bit activation quantization, significantly outperforming BiLLM’s 21.35 without activation quantization, underscoring its potential in compressing LLMs for real-world edge applications. Our code is available at: https://github.com/ZijianYY/DBellQuant.
Machine Learning, ICML
@@ 1 Introduction
In recent years, the rapid advancement of large language models (LLMs) demonstrate exceptional performance in a variety of complex tasks that involve natural language understanding and generation (achiam2023gpt; dubey2024llama). However, these models often comprise hundreds of billions of parameters, posing significant challenges for their deployment in real-world edge applications because of the substantial computational and memory requirements (e.g. a 70B model requires around 150GB GPU memory), resulting in huge operational costs and unacceptable inference latency.
@@
Figure 1: Performance on Wikitext2 dataset. DBellQuant outperforms weight-only quantization method under 8-bit activation setting.
@@
Quantization is a compelling route to LLM compression, and weight binarization is especially attractive for sharply reducing model size (wang2024bitnet; yu2024super). Recent PTQ methods– e.g., PB-LLM and BiLLM– mitigate binarization error with finer-grained treatment of salient weights (shang2023pb; huang2024billm). However, activation outliers remain a major obstacle to simultaneously quantizing activations (sun2024massive), undercutting memory and latency gains. Scale-redistribution techniques (xiao2023smoothquant; shao2023omniquant) and Hadamard-based transforms (ashkboos2024quarot; liu2024spinquant; sun2025flatquantflatnessmattersllm; lin2024duquantdistributingoutliersdual) help at higher weight precisions, but no PTQ weight binarization method concurrently achieves effective activation quantization. This is because pushing weights to 1-bit increases sensitivity to activation outliers, while existing fixes shift the difficulty back to the weights or require higher-bit activations for accuracy. Quantization-aware training (QAT) approaches such as BitNet can reach 1-bit weights with low-bit activations (wang2024bitnet), but at the cost of heavy retraining. Thus, there is a need for post-training weight binarization that preserves accuracy while also enabling activation quantization.
To this end, we begin by revisiting the distribution characteristics of activations and weights in LLMs (Fig. 2(a)). We observe that the unimodal nature of weight distributions leads to substantial quantization errors, particularly in the case of low 1-bit quantization. Ideally, a dual-bell-shaped weight distribution (Fig. 2(b)) can effectively reduce binarization errors. In addition, activation distribution suffer from outliers and thus large quantization errors, which demands a narrower distribution. This motivates a key question: Is it possible to transform weights into a dual-bell shape distribution while simultaneously addressing activation outliers to facilitate both activation quantization and weight binarization?
Building on these observations, we propose DBellQuant, a novel weight-activation quantization framework for efficient post-training quantization (PTQ). DBellQuant enables activation quantization while achieving near 1-bit weight compression with minimal accuracy loss. The core mechanism is a learnable, equivalence-preserving transformation that maps each layer’s weight distribution toward a dual-bell shape; its inverse is applied to the input activations to keep the computation unchanged. We analyze the feasibility of dual-bell transformations for binarization and propose a lightweight algorithm, Learnable Transformation for Dual-Bell (LTDB), which initializes the transform and optimizes it with a custom objective that drives weights to cluster around two modes, with early stopping to ensure stability. By doing so, weights originally exhibiting unimodal distributions– challenging for binarization– are mapped into near-symmetric dual-bell distributions (Fig. 2(b)), significantly reducing binarization error. Moreover, as discussed in Sec. 3.4, experiments show that applying the inverse transform to inputs consistently contracts activation ranges and suppresses outliers, making activations more amenable to quantization (Fig. 2(b)).
Experimental results demonstrate that the equivalent transformation strategy of DBellQuant significantly reduces the loss caused by weight binarization. For the first time under PTQ conditions, it achieves near 1-bit weight compression while simultaneously compressing activations to 6 bits or even 4bits. Across various LLMs and evaluation metrics, DBellQuant consistently achieves state-of-the-art results. For instance, on the Wikitext2 benchmark, we achieve a perplexity of 21.69 on LLaMA-2-7B using only 6-bit activations, even with 4-bit activations, we achieve a perplexity of 23.04, significantly surpassing the performance of BiLLM, which achieves a perplexity of 21.35 .
@@ 2 Related Work
The massive parameter size of LLMs poses significant challenges in terms of memory consumption and computational efficiency. Therefore, quantization is crucial to compress these models, reducing resource requirements while preserving performance for practical deployment. LLMs quantization have introduced a variety of innovative techniques to enhance efficiency while maintaining accuracy. Works like GPTQ (frantar2022gptq) and OBQ (frantar2023optimalbraincompressionframework) minimizes reconstruction error by adjusting the remaining unquantized parameters in the block to compensate for the accuracy loss caused by quantization. LLM.int8()(dettmers2022gpt3) and ZeroQuant(yao2022zeroquant) improve quantization accuracy by introducing additional grouping labels for customized quantization blocks. Other works like SmoothQuant(xiao2023smoothquant) and OmniQuant (shao2023omniquant) addresses activation outliers by redistributing scaling factors between weights and activations, migrating the quantization difficulty from activation to weights. Additionally, recent approaches leverage Hadamard transformations to suppress activation outliers (ashkboos2024quarot; liu2024spinquant; sun2025flatquantflatnessmattersllm), while incoherence processing and non-linear transformation have been proposed for effective low-bit quantization (chee2024quip2bitquantizationlarge; tseng2024quipbetterllmquantization; zhang2024magrweightmagnitudereduction). Collectively, these advancements demonstrate that quantization techniques can be scaled to multi-billion-parameter models, achieving substantial reductions in memory consumption and inference latency without compromising model performance.
@@
Binary quantization, an extreme low-bit quantization technique that reduces model weights and activations to binary values (e.g., -1 and +1 or 0 and 1), has gained significant attention for its ability to drastically cut memory usage and computational complexity, making it ideal for resource-constrained devices and efficient deployment of large-scale models. However, applying binary quantization to LLMs presents substantial challenges due to their sensitivity to precision loss, particularly in attention mechanisms and large embedding layers. BinaryBERT (bai2020binarybert) explored binary quantization for BERT, proposing selective preservation of critical weights in higher precision to mitigate performance degradation. In another direction, PB-LLM (shang2023pb) introduced a partially-binarized approach for LLMs, retaining a small fraction of salient weights in higher precision while binarizing the rest, enabling extreme low-bit quantization without sacrificing linguistic reasoning capabilities. Recent advancements include structural binarization techniques that leverage novel sparsity forms and standardized importance metrics to selectively binarize and sparsify LLM weights (dong2024stbllmbreaking1bitbarrier), as well as strategies like alternating refined binarization and column-group bitmap methods to effectively reduce quantization error and address column deviations (li2024arbllmalternatingrefinedbinarizations). These innovations collectively advance the feasibility of binary quantization for LLMs, pushing the boundaries of efficiency without compromising performance.
@@ 3.1 Binarization-Friendly Weight Redistribution
$\tiny\beta=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{W_{i,j}},\quad\boldsymbol{\widetilde{W}}=\text{Sign}(\boldsymbol{W}-\beta),\quad\alpha=\frac{1}{n}\sum_{i=1}^{n}|\boldsymbol{W_{i,j}}-\beta_{j}|$ (1)
$\tiny\text{Sign}(\boldsymbol{W_{i,j}})=\begin{cases}+1,&\text{if }\boldsymbol{W_{i,j}}>0,\\ -1,&\text{if }\boldsymbol{W_{i,j}}\leq 0,\end{cases},\quad\boldsymbol{W_{deq}}=\boldsymbol{\widetilde{W}}\cdot\alpha+\beta$ (2)
where $\beta$ is the shifting factor and $\alpha$ is the scaling factor for binarization. Previous studies (huang2024slim) have shown that neural network weights exhibit structured distributions along the channel dimension, with certain channels being more salient. The overall weight distribution typically follows a quasi-Gaussian pattern (dettmers2022llmint88bitmatrixmultiplication), as does the channel-wise distribution (Fig. 6). Binarizing such weight matrices introduces significant quantization errors, which can severely degrade model performance.
In LLM binarization, a dual-bell distribution is theoretically more advantageous than a single-bell distribution due to its natural separation into two distinct clusters, which aligns well with binary quantization levels (e.g., -1 and 1), thereby minimizing quantization error. In contrast, single-bell distributions, concentrated around a single peak, often cause significant overlap when mapped to binary values, reducing representation accuracy (see Appendix A.20 for detailed analysis). However, LLM weight distributions typically exhibit single-bell characteristics, and conventional PTQ methods fail to effectively transform them for binarization. While QAT can reshape weight distributions into a dual-bell form through its learning objectives (wang2023bitnetscaling1bittransformers), it requires substantial computational resources and prolonged training. To address this, we propose a more efficient PTQ method that rapidly converts single-bell distributions into dual-bell ones, optimizing binary quantization without the need for resource-intensive retraining.
@@ 3.2 Learnable Transformation for Dual-Bell Quantization
Learnable Transformation with Auxiliary Matrix
@@
where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are parameters of the double-bell distribution. More detailed proof is shown in Appendix. A.19.
@@ 3.3 Dual-Transformation Optimizing Objectives
Our learning objective is to encourage all weight values to move closer to the two mean centers calculated by Eq. 2, ultimately forming a double-bell distribution with these two values as its peaks. During the binarization process, we denote these two points as $m_{1}$ and $m_{2}$ respectively. The simple way to set loss function is as follows:
@@
$\tiny\mathcal{L}_{\text{DTNP}}=\frac{\lambda_{DTNP}}{n}\sum_{i=1}^{n}\begin{cases}\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|}{\left|\text{m}_{1,i}\right|},&\text{if }\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|<\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|\\[10.0pt] \frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|}{\left|\text{m}_{2,i}\right|},&\text{otherwise.}\end{cases}$ (6)
@@ 3.4 Impact of the Inverse of Learnable Transformation Matrix on Activation Smoothing
Driven by the DTNP objective, the transformation matrix $T$ encourages weight values to cluster around the two binarization centers defined in Eq. 2. Given that LLM weights typically follow a quasi-Gaussian distribution concentrated near zero (dettmers2022llmint88bitmatrixmultiplication), $T$ tends to learn values greater than 1 to amplify these small magnitudes towards the binarization targets. Crucially, this mechanism inherently benefits activation quantization. Our activation-aware initialization (Eq. 4) establishes a positive correlation between the activation magnitude $\max(|X_{j}|)$ and the transformation scale $T_{j}$. Consequently, channels exhibiting significant activation outliers are assigned larger $T_{j}$ values. When the inverse transformation $T^{-1}$ is applied to the input $X$, these outlier-heavy channels undergo stronger suppression (multiplication by a smaller $1/T_{j}$ factor) compared to normal channels. Empirically, as visualized in Fig. 5, we observe that less than 5% of values in $T$ are below 1, implying that $T^{-1}$ acts as a global compressor with channel-specific intensity. This effectively narrows the dynamic range of activations and mitigates the kurtosis caused by outliers. We provide a quantitative validation of this smoothing effect in Appendix A.17, where metrics including Z-score and Relative Deviation Error demonstrate a substantial reduction in distribution variance and outlier impact, making the transformed activations significantly more amenable to low-bit quantization.
@@ 3.5 Algorithm
The process of Learnable Transformation for Dual-Bell Transformation Algorithm is shown in Algorithm. 1, which can be seen in Appendix. A.18 and can adjust the full-precision weight matrix $W$ using an activation-aware initialization transformation matrix $T$ over $N$ epochs. The algorithm iteratively minimizes dual-target loss functions to guide $W$ toward a dual-bell distribution while employing an early stopping mechanism based on the DTMD.